# Flattening Hierarchies with Policy Bootstrapping

### v1

@@ C.2 RIS and SAW derivations
Unlike HIQL, both RIS and SAW seek to learn a unified flat policy, and therefore we choose a policy posterior $\pi_{\theta}(\tau)$ that factors as
$\pi_{\theta}(\tau\mid g)=p(s_{0})\prod^{\infty}_{t=0}p(s_{t+1}\mid s_{t},a_{t})\pi_{\theta}(a_{t}\mid s_{t},g).$
We also introduce a variational posterior $q$ which factors over a sequence of waypoints $\{w\}=\{w_{0},w_{1},\ldots\}$ as
$q(\{w\}\mid g)=p(s_{0})\prod^{\infty}_{t=0}p(s_{t+1}\mid s_{t},a_{t})\pi^{\mathrm{sub}}(a_{t}\mid s_{t},w_{t})q(w_{t}\mid s_{t},g),$
@@
where $\pi^{\mathrm{sub}}$ is a target subpolicy and is treated as fixed with respect to the parameters of the posteriors. Using these definitions, we define the evidence lower bound (ELBO) on the likelihood of subgoal optimality $p_{\pi}(U=1)$ for policy $\pi$
[…]
$=\log\mathbb{E}_{\pi_{\theta}(\tau\mid g),q(\{w\}\mid g),p(g)}\left[\frac{p(U=1,\tau,\{w\}\mid g)}{\pi_{\theta}(\tau\mid g)q(\{w\}\mid g)}\right]$
$\geq\mathbb{E}_{\pi_{\theta}(\tau\mid g),q(\{w\}\mid g),p(g)}\log\left[\frac{p(U=1,\tau,\{w\}\mid g)}{\pi_{\theta}(\tau\mid g)q(\{w\}\mid g)}\right]=\mathcal{J}(\pi_{\theta},q).$
@@
$\mathbb{E}_{\pi_{\theta}(\tau\mid g),\,q(\{w\}\mid g),\,p(g)}\left[\beta\sum_{t=0}^{\infty}\gamma^{t}A(s_{t},w_{t},g)+\sum_{t=0}^{\infty}\log\left[\frac{\pi^{\ell}(a_{t}\mid s_{t},w_{t})\,\pi^{h}(w_{t}\mid s_{t},g)}{\pi_{\theta}(a_{t}\mid s_{t},g)\,q(w_{t}\mid s_{t},g)}\right]\right]$
Once again, we express the discounted sum over time as an expectation over the discounted stationary state distribution $\mu(s)$ and omit the conditionals in the expectation over $\pi_{\theta}(a)$ and $q(w)$ for brevity. Simplifying gives us
$\mathbb{E}_{\mu(s),\pi_{\theta}(a),q(w),\,p(g)}\left[\beta A(s,w,g)+\log\left[\frac{\pi^{h}(w\mid s,g)}{q(w\mid s,g)}\right]+\log\left[\frac{\pi^{\ell}(a\mid s,w)}{\pi_{\theta}(a\mid s,g)}\right]\right]$
$=\mathbb{E}_{\mu(s),q(w),\,p(g)}\left[\beta A(s,w,g)\right]-\mathbb{E}_{\mu(s),\,p(g)}\left[D_{\mathrm{KL}}\left(q\left(w\mid s,g\right)\|\pi^{h}\left(w\mid s,g\right)\right)\right]$
$\qquad+\mathbb{E}_{\mu(s),q(w),\,p(g)}\left[D_{\mathrm{KL}}\left(\pi_{\theta}\left(a\mid s,g\right)\|\pi^{\ell}\left(a\mid s,w\right)\right)\right]$
RIS: RIS partitions this objective into two parts and optimizes them separately, where the subgoal generator is trained according to the loss
$\mathcal{J}^{h}(q)=\mathbb{E}_{\mu(s),q(w),\,p(g)}[A(s,w,g)]-\mathbb{E}_{\mu(s),\,p(g)}\left[D_{\mathrm{KL}}(q(w\mid s,g)\|\pi^{h}(w\mid s,g))\right],$
which is identical to the objective for the HIQL high-level policy $\pi_{\psi}$ and yields the same AWR-like objective over subgoals $\exp(A(s,w,g))\log q(w\mid s,g)$. We then incorporate the remaining KL divergence term into the objective for the flat policy posterior $\pi_{\theta}$
$\mathcal{J}(\theta)=\mathbb{E}_{\mu(s),q(w),\,p(g)}[D_{\mathrm{KL}}(\pi_{\theta}(a\mid s,g)\|\pi^{\ell}(a\mid s,w))],$
which is an expectation over subgoals drawn from the (simultaneously learned) subgoal generator.
SAW: Instead of directly learning $q$, we use Bayes’ rule and our earlier definition of $p(U=1\mid s,w,g)$ to directly approximate the posterior distribution over subgoals $p(w\mid s,g,U=1)$, where
@@
Although the proportionality constant in the first line is the $p_{\pi}(U=1)$, which is the subject of our optimization, we note that approximating the expectation over subgoals $w$ corresponds to the expectation (E) step in a standard expectation-maximization (EM) procedure [1]. Because we are only seeking to fit the shape of the optimal variational posterior over subgoals for the purposes of approximating the expectation over $q$, and not maximizing $p_{\pi}(U=1)$ (the M step), we can treat $p_{\pi}(U=1)$ as constant with respect to $q$ to get
$\mathcal{J}(\theta)$ $=\mathbb{E}_{\mu(s),q(w),\,p(g)}[D_{\mathrm{KL}}(\pi_{\theta}(a\mid s,g)\|\pi^{\ell}(a\mid s,w))]$
$=\int\mu(s)\,p(g)\,q(w\mid s)[D_{\mathrm{KL}}(\pi_{\theta}(a\mid s,g)\|\pi^{\ell}(a\mid s,w))]dw\,dg\,ds$
@@
which yields our subgoal advantage-weighted bootstrapping term in Equation 6.
Appendix D Offline GCRL Baseline Algorithms
In this section, we briefly review the baseline algorithms referenced in Table 1. For more thorough implementation details, as well as goal-sampling distributions, interested readers may refer to Appendix C of Park et al. [31] as well as the original works.
@@ Appendix D Offline GCRL Baseline Algorithms
Goal-conditioned implicit {Q, V}-learning (GCIQL & GCIVL): GCIQL is a goal-conditioned variant of implicit Q-learning [20], which performs policy iteration with an expectile regression to avoid querying the learned $Q$-value function for out-of-distribution actions. Park et al. [33] introduced a $V$-only variant that directly regresses towards high-value transitions [Equation 2], using $r(s,g)+\gamma\bar{V}\left(s^{\prime},g\right)$ as an estimator of $Q(s,a,g)$. Since it does not learn $Q$-values and therefore cannot marginalize over non-causal factors, it is optimistically biased in stochastic environments.
@@
Although both baselines are value learning methods that can be used with multiple policy extraction objectives (including our own, which uses GCIVL), the OGBench implementations are paired with following objectives: Deep Deterministic Policy Gradient with a behavior cloning penalty term [11, DDPG+BC] for GCIQL, and AWR [Equation 3] for GCIVL.
@@
Quasimetric RL (QRL): QRL [47] is a non-traditional value learning algorithm that uses Interval Quasimetric Embeddings [46, IQE] to enforce quasimetric properties (namely, the triangle inequality and identity of indiscernibles) of the goal-conditioned value function on distances between representations. It uses a constrained “maximal spreading” objective to estimate the shortest paths between states, then learns a one-step dynamics model combined with DDPG+BC to extract a policy from the learned representations.
@@
Contrastive RL (CRL): CRL [10] is a representation learning algorithm which uses contrastive learning to enforce that the inner product between the learned representations of a state-action pair and the goal state corresponds to the discounted future state occupancy measure of the goal state, which is estimated directly from data using Monte Carlo sampling. CRL then performs one-step policy improvement by choosing actions that maximize the future occupancy of the desired goal state.
@@
Hierarchical implicit Q-learning (HIQL): HIQL [33] is a policy extraction method that learns two levels of hierarchical policy from the same goal-conditioned value function. The low-level policy $\pi^{\ell}$ is trained using standard AWR, and the high-level policy $\pi^{h}$ is trained using an action-free, multi-step variant of AWR that treats (latent) subgoal states as “actions.”
@@
Reinforcement learning with imagined subgoals (RIS): RIS [6] is a policy extraction method originally designed for the online GCRL setting, which learns a subgoal generator and a flat, goal-conditioned policy. Unlike SAW, RIS uses a fixed coefficient on the KL term, instead learning a subgoal generator and bootstrapping directly on a target policy (parameterized by an exponential moving average of online policy parameters rather than a separately learned subpolicy) conditioned on “imagined” subgoals. It also incorporates a value-based policy learning objective similar to our approach, but learns a $Q$-function and differentiates directly through the policy with DDPG.
@@
To modify RIS for the offline setting in our implementation for Figure 3, we fixed the coefficient on the KL term to $\beta=3.0$, trained a subgoal generator identical to the one in HIQL, and replaced the dataset subgoals in SAW with “imagined” subgoals. Otherwise, for fairness of comparison, our offline RIS implementation used the same hyperparameters and architectures as SAW, including a separate target subpolicy network instead of a soft copy of the online policy, subgoals at a fixed distance instead of at midpoints, and AWR instead of DDPG+BC for the policy extraction objective, which we found to perform better in locomotion environments.
@@ Appendix F Implementation Details
Target policy: While Chane-Sane et al. [6] use a exponential moving average (EMA) of the online policy parameters $\theta$ as the target policy prior $\pi_{\overline{\theta}}$, we instead simply train a smaller policy network parameterized separately by $\psi$ on (sub)goals sampled from $k$ steps into the future, where $k$ is a hyperparameter. We find that this leads to faster training and convergence, albeit with a small increase in computational complexity.
@@
Architecture: During our experiments, we observed that the choice of network architecture for both the value function and policy networks had a significant impact on performance in several environments. Instead of taking in the raw concatenated state and goal inputs, HIQL prepends a subgoal representation module consisting of an additional three-layer MLP followed by a bottleneck layer of dimension 10 and a length-normalizing layer that projects state-(sub)goal representations to the unit hypersphere. The value and low-level policy networks receive this representation in place of the goal information, as well as the raw (in state-based environments) or encoded (in pixel-based environments) state information. We found that simply adding these additional layers (separately) to the value and actor network encoders significantly boosted performance in state-based locomotion tasks, with modifications to the former improving training stability in pointmaze and modifications to the latter being critical for good performance in the antmaze and humanoidmaze environments.
@@
While we did not perform comprehensive architectural ablations due to computational limitations, we note that the desirable properties of the unit hypersphere as a representation space are well-studied in contrastive learning [45] and preliminary work by Wang et al. [44] has explored the benefits of scaling network depth for GCRL (albeit with negative results for the offline setting). Further studying the properties of representations emerging from these architectural choices may inform future work in representation learning for offline GCRL.

### v3

@@ B.2 RIS and SAW derivations
Unlike HIQL, RIS and SAW both seek to learn a unified flat policy, and therefore we choose a policy posterior $q^{f}(\tau)$ that factors as
$q^{f}(\tau\mid g)=p(s_{0})\prod^{\infty}_{t=0}p(s_{t+1}\mid s_{t},a_{t})q^{f}(a_{t}\mid s_{t},g).$
For RIS, which learns a generative subgoal policy identical to that of HIQL, we also use the same distribution $q^{h}$ which factors over a sequence of waypoints $\{w\}=\{w_{0},w_{1},\ldots\}$ as
$q^{h}(\{w\}\mid g)=p(s_{0})\prod^{\infty}_{t=0}p(s_{t+1}\mid s_{t},a_{t})\pi^{\mathrm{sub}}(a_{t}\mid s_{t},w_{t})q^{h}(w_{t}\mid s_{t},g).$
@@
Rather than jointly learning low- and high-level policies, RIS bootstraps from a target subpolicy $\pi^{\mathrm{sub}}$, which is treated as fixed. Using these definitions, we define the evidence lower bound (ELBO) on the likelihood of subgoal optimality $p_{\pi}(U=1)$ for policy $\pi$
$\log p_{\pi}(U=1)$ $=\log\int p(g)p(U=1,\tau,\{w\}\mid g)d\tau d\{w\}dg$
$=\log\int p(g)q^{f}(\tau\mid g)q^{h}(\{w\}\mid g)\frac{p(U=1,\tau,\{w\}\mid g)}{q^{f}(\tau\mid g)q^{h}(\{w\}\mid g)}d\{w\}\,d\tau\,dg$
$=\log\mathbb{E}_{q^{f}(\tau\mid g),q^{h}(\{w\}\mid g),p(g)}\left[\frac{p(U=1,\tau,\{w\}\mid g)}{q^{f}(\tau\mid g)q^{h}(\{w\}\mid g)}\right]$
$\geq\mathbb{E}_{q^{f}(\tau\mid g),q^{h}(\{w\}\mid g),p(g)}\log\left[\frac{p(U=1,\tau,\{w\}\mid g)}{q^{f}(\tau\mid g)q^{h}(\{w\}\mid g)}\right]=\mathcal{J}(q,\pi).$
@@
$\mathbb{E}_{q^{f}(\tau\mid g),\,q^{h}(\{w\}\mid g),\,p(g)}\left[\beta\sum_{t=0}^{\infty}\gamma^{t}A(s_{t},w_{t},g)+\sum_{t=0}^{\infty}\log\left[\frac{\pi^{\ell}(a_{t}\mid s_{t},w_{t})\,\pi^{h}(w_{t}\mid s_{t},g)}{q^{f}(a_{t}\mid s_{t},g)\,q^{h}(w_{t}\mid s_{t},g)}\right]\right].$
Once again, we express the discounted sum over time as an expectation over the discounted stationary state distribution $\mu(s)$ and omit the conditionals in the expectation over $q^{f}(a)$ and $q^{h}(w)$ for brevity. Simplifying gives us
$\mathbb{E}_{\mu(s),q^{f}(a),q^{h}(w),\,p(g)}\left[\beta A(s,w,g)+\log\left[\frac{\pi^{h}(w\mid s,g)}{q^{h}(w\mid s,g)}\right]+\log\left[\frac{\pi^{\ell}(a\mid s,w)}{q^{f}(a\mid s,g)}\right]\right]$
$=\mathbb{E}_{\mu(s),q^{h}(w),\,p(g)}\left[\beta A(s,w,g)\right]-\mathbb{E}_{\mu(s),\,p(g)}\left[D_{\mathrm{KL}}\left(q^{h}(w\mid s,g)\|\pi^{h}\left(w\mid s,g\right)\right)\right]$
$\qquad+\mathbb{E}_{\mu(s),q^{h}(w),\,p(g)}\left[D_{\mathrm{KL}}\left(q^{f}\left(a\mid s,g\right)\|\pi^{\ell}\left(a\mid s,w\right)\right)\right].$
RIS: RIS partitions this objective into two parts and optimizes them separately. The first line is identical to the HIQL high-level policy objective and yields the same AWR-like objective for a parametric generative subgoal policy $\pi^{h}_{\psi}$
$\mathcal{J}^{h}_{\mathrm{RIS}}(\psi)$ $=\mathbb{E}_{\mu(s),q^{h}(w),p(g)}\left[\exp(A(s,w,g))\log\pi^{h}_{\psi}(w\mid s,g)\right].$
The remaining KL divergence term seeks to minimize the divergence between the flat goal-conditioned policy posterior $q^{f}$ and the subgoal-conditioned policy $\pi^{\ell}(a\mid s,w)$, which is fit separately with the target subpolicy $\pi^{\mathrm{sub}}(a\mid s,w)$ from earlier. Importantly, because $\pi^{\mathrm{sub}}(a\mid s,w)$ is (presumably) a parametric distribution, we can directly optimize the KL divergence in the space of policies with another parametric policy $\pi_{\theta}(a\mid s,g)$, yielding the final bootstrapping objective
$\mathcal{J}^{f}_{\mathrm{RIS}}(\theta)=\mathbb{E}_{\mu(s),q^{h}(w),\,p(g)}[D_{\mathrm{KL}}(\pi_{\theta}(a\mid s,g)\|\pi^{\mathrm{sub}}(a\mid s,w))].$
This divergence is minimized in expectation over optimal subgoals $q^{h}(w\mid s,g)$, which in turn is approximated by the parametric subgoal policy $\pi^{h}_{\psi}(w\mid s,g)$.
SAW: Instead of fitting a generative model $q^{h}(w\mid s,g)$ over the potentially high-dimensional space of subgoals, we can use a simple application of Bayes’ rule to directly approximate the expectation over the optimality-conditioned distribution of subgoals $p(w\mid s,g,U=1)$, where
@@
Although the proportionality constant in the first line is the $p_{\pi}(U=1)$, which is the subject of our optimization, we note that approximating the expectation over subgoals $w$ corresponds to the expectation step in a standard expectation-maximization (EM) procedure [abdolmaleki_maximum_2018]. Because we are only seeking to fit the shape of the optimal posterior over subgoals for the purposes of approximating the expectation over $q^{h}$, and not maximizing $p_{\pi}(U=1)$ (the M step), we can treat $p_{\pi}(U=1)$ as constant with respect to $q^{h}$ to get
$\mathcal{J}_{\mathrm{SAW}}(\theta)$ $=\mathbb{E}_{\mu(s),q^{h}(w),\,p(g)}[D_{\mathrm{KL}}(\pi_{\theta}(a\mid s,g)\|\pi^{\ell}(a\mid s,w))]$
$=\int\mu(s)\,p(g)\,q^{h}(w\mid s)[D_{\mathrm{KL}}(\pi_{\theta}(a\mid s,g)\|\pi^{\ell}(a\mid s,w))]dw\,dg\,ds$
@@
giving us the final subgoal advantage-weighted bootstrapping term in Equation 8.
Appendix C Offline GCRL Baseline Algorithms
In this section, we briefly review the baseline algorithms referenced in Table 1. For more thorough implementation details, as well as goal-sampling distributions, interested readers may refer to Appendix C of park_ogbench_2024 as well as the original works.
@@ Appendix C Offline GCRL Baseline Algorithms
Goal-conditioned implicit {Q, V}-learning (GCIQL & GCIVL): GCIQL is a goal-conditioned variant of implicit Q-learning [kostrikov_offline_2021], which performs policy iteration with an expectile regression to avoid querying the learned Q-value function for out-of-distribution actions. park_hiql_2023 introduced a $V$-only variant that directly regresses towards high-value transitions [Equation 2], using $r(s,g)+\gamma\bar{V}\left(s^{\prime},g\right)$ as an estimator of $Q(s,a,g)$. Since it does not learn Q-values and therefore cannot marginalize over non-causal factors, it is optimistically biased in stochastic environments.
@@
Although both baselines are value learning methods that can be used with multiple policy extraction objectives (including SAW, which extracts a policy from a value function learned with GCIVL), the OGBench implementations are paired with following objectives: Deep Deterministic Policy Gradient with a behavior cloning penalty term [fujimoto_minimalist_2021, DDPG+BC] for GCIQL, and AWR [Equation 3] for GCIVL.
@@
Quasimetric RL (QRL): QRL [wang_optimal_2023] is a non-traditional value learning algorithm that uses Interval Quasimetric Embeddings [wang_improved_2022, IQE] to enforce quasimetric properties (namely, the triangle inequality and identity of indiscernibles) of the goal-conditioned value function. It relies on the Lagrangian dual form of a constrained “maximal spreading” objective to estimate the shortest paths between states, then learns a one-step dynamics model combined with DDPG+BC to extract a policy from the learned representations.
@@
Contrastive RL (CRL): CRL [eysenbach_contrastive_2022] is a representation learning algorithm which uses contrastive learning to enforce that the inner product between the learned representations of a state-action pair and the goal state corresponds to the discounted future state occupancy measure of the goal state, which is estimated directly from data using Monte Carlo sampling. CRL then performs one-step policy improvement by choosing actions that maximize the future occupancy of the desired goal state.
@@
Hierarchical implicit Q-learning (HIQL): HIQL [park_hiql_2023] is a policy extraction method that learns two levels of hierarchical policy from the same goal-conditioned value function. The low-level policy $\pi^{\ell}$ is trained using standard AWR, and the high-level policy $\pi^{h}$ is trained using an action-free, multi-step variant of AWR that treats (latent) subgoal states as “actions.”
@@
Reinforcement learning with imagined subgoals (RIS): RIS [chane-sane_goal-conditioned_2021] is a policy extraction method originally designed for the online GCRL setting, which learns a subgoal generator and a flat, goal-conditioned policy. Unlike SAW, RIS uses a fixed coefficient on the KL term, instead learning a subgoal generator and bootstrapping directly on a target policy (parameterized by an exponential moving average of online policy parameters rather than a separately learned subpolicy) conditioned on “imagined” subgoals. It also incorporates a value-based policy learning objective similar to our approach, but learns a $Q$-function and differentiates directly through the policy with DDPG.
@@
Offline RIS (RISoff): To modify RIS for the offline setting in our implementation for Figure 3, we fixed the coefficient on the KL term to $\beta=100.0$ for navigation and $\beta=3.0$ for manipulation tasks, trained a subgoal generator identical to the one in HIQL, and replaced the dataset subgoals in SAW with “imagined” subgoals. Otherwise, for fairness of comparison, our offline RIS implementation used the same hyperparameters and architectures as SAW, including a separate target subpolicy network instead of a soft copy of the online policy, subgoals at a fixed distance instead of at midpoints, and AWR instead of DDPG+BC for the policy extraction objective, which we found to perform better in locomotion environments.
Appendix D Links to Planning Invariance
As an aside, we note that the discussions in this paper are closely related to the recently introduced concept of planning invariance [myers_horizon_2024], which describes a policy that takes similar actions when directed towards a goal as when directed towards an intermediate waypoint en route to that goal. In fact, we can say that subgoal-conditioned HRL methods achieve a form of planning invariance by construction, since they simply use the actions yielded by waypoint-conditioned policies to reach further goals. By minimizing the divergence between the full goal-conditioned policy and an associated subgoal-conditioned policy, both SAW and RIS can also be seen as implicitly enforcing planning invariance.
@@ Appendix F Implementation Details
F.1 Simplified advantage estimates
GCRL typically uses a sparse reward formulation where the reward is always constant (typically 1-1): in other words, the goal-conditioned reward function is always r(s,g)=1r(s,g)=-1 unless the current state is within some small distance $\epsilon$ from the goal $g$, where $r(s_{g},g)=0$. While the full advantage estimate would be $A\left(s_{t},w,g\right)=-V\left(s_{t},g\right)+\gamma^{k}V(w,g)+\sum_{t^{\prime}=t}^{k-1}r\left(s_{t^{\prime}},g\right)$, the sum over rewards should almost always be a constant when combined with hindsight goal relabeling, unless the goal was already reached at some point between time $t$ and t+k1t+k-1. Finally, the discount $\gamma^{k}$ is always a constant (since we fix $k$ as a hyperparameter), and can therefore be roughly folded into the inverse temperature parameter $\beta$ in the subgoal advantage $\exp(\beta A(s,w,g))$.
While this is not a completely faithful approximation, we found in our experiments that it yielded similar results and even improved performance in some cases by reducing the absolute magnitude of the advantage term in the exponent and therefore the “sharpness” of the resulting distribution, compared to the full expression that applies the discount factor $\gamma^{k}$ to $V(w,g)$.
F.2 Goal sampling
Following the standard hyperparameter settings in OGBench, we use separate goal-sampling distributions to train the value networks and actors. The value functions are trained with a mixture distribution $p_{\mathrm{mix}}(g\mid s)$ where 20% of goals are simply set to the current state $s_{t}$, 50% are sampled from future states $s_{t+k}$ in the same trajectory, using geometric sampling where $k\sim\mathrm{Geom}(1-\gamma)$, and 30% are sampled uniformly at random from the entire dataset $\mathcal{D}$. On the other hand, the policy is trained with future states sampled uniformly from the same trajectory.
F.3 Target policy
@@
While chane-sane_goal-conditioned_2021 use an exponential moving average (EMA) of the online policy parameters $\theta$ as the target policy prior $\pi_{\overline{\theta}}$, we instead simply train a smaller policy network parameterized separately by $\psi$ on (sub)goals sampled from $k$ steps into the future, where $k$ is a hyperparameter. We find that this leads to faster convergence, albeit with a small increase in computational complexity.
F.4 Architectures
Apart from the goal encoder discussed below, all networks use 3-layer multilayer perceptrons (MLPs) of dimension 512 with layer normalization and Gaussian error linear units (gelus). For environments with pixel-based observations, we additionally use a small Impala encoder [espeholt_impala_2018] with a single Resnet block [he_deep_2016]. For value learning, we use an ensemble of two networks for both the online and target networks with double-Q learning as in the GCIVL implementation. Importantly, only the full policy and not the target subpolicy actor $\pi^{\mathrm{sub}}$ uses the goal encoder.
@@ F.4 Architectures
Goal encoder: During our experiments, we observed that the choice of network architecture for both the value function and policy networks had a significant impact on performance in several environments. Instead of taking in the raw concatenated state and goal inputs, HIQL processes the concatenated current observation and goal states with a representation module consisting of an additional three-layer MLP for state-based environments, followed by a length-normalized linear projection to a 10-dimensional bottleneck representation. This is appended to another Impala encoder in visual environments. The value and low-level policy networks receive this representation in place of the goal information, as well as the raw (in state-based environments) or encoded (in pixel-based environments, using a separate Impala encoder) state information. While SAW has no need for a subgoal representation module, we found that simply adding these additional layers (separately) to the value and actor network encoders significantly boosted performance in state-based locomotion tasks, with modifications to the former improving training stability in pointmaze and modifications to the latter being critical for good performance in the antmaze and humanoidmaze environments.
@@
While we did not perform comprehensive architectural ablations due to computational limitations, we note that the desirable properties of the unit hypersphere as a representation space are well-studied in contrastive learning [wang_understanding_2020] and preliminary work by wang_1000_2025 has explored the benefits of scaling network depth for GCRL (albeit with negative results for the offline setting). Further studying the properties of representations emerging from these architectural choices may inform future work in representation learning for offline GCRL.